How many grams of NaOH will be required to prepare 1 liter of a stock solution of 40% NaOH
1
Expert's answer
2020-08-28T06:26:13-0400

w=m(NaOH)/(m(NaOH)+m(H2O) 0,4=x/(x+1000) x=667 g